OpenStudy (anonymous):

(1/x) + (1/y) = (1/f). What is x(y)? Show me how to get the answer as well please =)

OpenStudy (anonymous):

Guess and check is the best method for this, because you can't solve for three variables in one equation. 0 is one answer that works.

OpenStudy (anonymous):

I don't really need the answer, but more like an equation. There are 10 options which use the variables f, fy, and y, all in different forms.

OpenStudy (phi):

You could add 1/x + 1/y to get (x+y)/xy and then solve for xy

OpenStudy (anonymous):

one answer is \[xy=xf+yf\]

OpenStudy (anonymous):

Isolate the 1/x term, then invert both sides: \[1/x = 1/f - 1/y\] \[x = 1/(1/f - 1/y)\]